OpenStudy (anonymous):

can you guys help me with one more its Evaluate f(x) = x^ 2 + 1 for f(10). a.2√2 b.101 c.21

OpenStudy (***[isuru]***):

Hi, substitute 10 instead of "x" in f(x) f(x) = x^2 +1 = 10^2 + 1 = 101 hope this will help ya!!

OpenStudy (anonymous):

These are harder can you help with this one too? If f(x) = x^ 2 + 1, find f(a + 1). a.a^2 + 2 b.a ^2+ a + 1 c.a ^2+ 2a + 2

OpenStudy (***[isuru]***):

hi, use the same method substitute (a + 1) instead of "x" f(x) = x^2 + 1 = (a + 1)^2 + 1 = a^2 +2a +1 +1 = a^2 + 2a +2

OpenStudy (anonymous):

thanks how about Given the function f(x) = 3x + 1, evaluate f(a + 1). a.3a + 1 b.3a + 2 c.3a + 4

OpenStudy (***[isuru]***):

use the same old method here lol f(a+1) = 3(a + 1) +1 now simplify it 3a +1 +1 3a +2

OpenStudy (anonymous):

ok last one If f(x) = 3x + 1, then f(a + h) - f(a) = a.h b.3h c.3h + 2

OpenStudy (agent0smith):

At least try it yourself first, magan... follow the examples @***[ISURU]*** showed.